Question:- Like everything in Minecraft game, an Astronomer started thinking what will happen if our Sun would have been a square of similar luminosity.


a. If the Sun was a cube, what would be its surface temperature?


b. In what interval, does its magnitude lie as seen from the Earth?

Hint: The volume for Cube Sun remains the same as usual.


1
Expert's answer
2022-03-13T18:47:19-0400

The current temperature of the Sun surface is "T_1=5772\:\rm K". The temperature is related with luminosity by an equation

"L=\sigma T^4A"

Hence

"T\sim\sqrt[4]{L}"

(a) When the luminosity of Sun would have been a squared, then

"T_2=T_1\sqrt[4]{2}=5772\:\rm K*\sqrt[4]{2}=6864\:\rm K"

(b) The current visual brightness of the Sun "m_1=-26.74^m". When the luminosity of Sun would have been a squared, then

"m_2=m_1-2.512\lg 2\\
=-26.74^m-0.756^m
=-27.50^m"